### v5

# Dual-Spaces Invariance as a Universal Criterion for Identifying Multifractal Critical States

Abstract
In Anderson localization physics, eigenstates of disordered quantum systems are commonly classified as extended, localized, or critical, distinguished by their distinct spatial structures in real space. While critical states are known to exhibit multifractal characteristics, a precise and operational criterion for characterizing critical states remains an open challenge. In this work, we address this challenge by revisiting criticality from a dual-spaces perspective that treats position and momentum representations on equal footing. Building on the Liu–Xia criterion, which characterizes critical states by the simultaneous vanishing of Lyapunov exponents ($\gamma=\gamma_{m}=0$) in both spaces, we show that this dual-spaces characterization captures an essential feature of critical states that is not limited to Lyapunov exponents. In particular, through numerical simulations, we demonstrate that the inverse participation ratio exhibits closely related scaling behavior in position and momentum space for critical states. This position-momentum correspondence clearly distinguishes critical states from extended or localized ones, which instead display a pronounced asymmetry between the two representations. Our results establish a robust and universal framework for precisely characterizing multifractal critical states in disordered quantum systems, and provide practical guidance for their identification in current quantum simulation platforms.
I Introduction
Anderson localization photonic ; Critical0 is a fundamental manifestation of wave interference in disordered media, leading to the suppression of diffusive transport Liu-self-duality ; Li1 ; Li2 . Within such systems, single-particle eigenstates are commonly classified as extended (metallic), localized (insulating), or critical. Critical states–also referred to as multifractal states–occupy an intermediate regime between extended and localized phases You ; g1 ; zhou . Unlike extended states, which exhibit nearly uniform probability distributions, or localized states, which decay exponentially, critical states typically display power-law spatial decay Biddle ; Zhang ; q2 . Their hallmark features include multifractality, characterized by a continuous spectrum of anomalous scaling exponents q1 ; q4 ; gg10 ; gg11 , as well as scale invariance and statistical self-similarity arising from the competition between localization and delocalization gg12 ; q5 ; q6 . In paradigmatic settings such as the three-dimensional Anderson model, critical states emerge at mobility edges separating extended and localized spectral regions Yao1 ; Yao2 .
Quasiperiodic systems cai ; Slager ; SSHQ ; gg3 ; Ayan1 ; Ayan2 provide an alternative platform for realizing critical states, even in the absence of true randomnes. Recent progress in critical states has uncovered a variety of nontrivial localization phenomena, including anomalous mobility edges Liu-Anomalous , renormalization-group descriptions RG1 ; RG2 , recursion-defined real spectra Liu-Real1 ; Liu-Real2 , and coupled-chain-induced criticality twochain . In previous theoretical descriptions Jiang1 ; Jiang2 ; gg5 ; gg6 , the asymptotic localization properties of single-particle eigenstates were typically characterized by a single real-space Lyapunov exponent $\gamma$. For exponentially localized states, the wave-function amplitudes decay as $|\psi_{n}|\propto e^{-\gamma|n-n_{0}|}$, where $n_{0}$ denotes the localization center and $\gamma>0$ quantifies the inverse localization length gg9 ; Duncan . A finite Lyapunov exponent $\gamma>0$ signals exponential localization, whereas $\gamma=0$ indicates the absence of exponential decay. It is important to emphasize, however, that a vanishing Lyapunov exponent $\gamma=0$ in real space does not uniquely characterize extended states: both truly extended states and critical (multifractal) states fall into the non-exponentially localized class. Consequently, a single-space description based solely on the real-space Lyapunov exponent suffers from intrinsic limitations when applied to critical states Liu-Anomalous .
Figure 1: (Color online) Schematic illustration of typical eigenstates in position and momentum space. Panels (a), (c), and (e) show extended, critical, and localized states in position space, respectively, while panels (b), (d), and (f) show their corresponding momentum-space distributions. In contrast to conventional approaches that emphasize real-space multifractality, a dual-space perspective highlights that critical states are distinguished by the absence of exponential localization in both representations.
A useful conceptual perspective is offered by Bohr’s complementarity principle MQM . Position (real) and momentum spaces constitute a pair of Fourier-conjugate representations: while position space captures the spatial distribution of particles, momentum space encodes their reciprocal structure. Extended and localized states typically display strongly asymmetric localization properties in these two representations. This observation naturally raises the question of whether critical states, which are self-similar by construction, may exhibit a more symmetric behavior under position-momentum transformation.
Motivated by this consideration, Liu and Xia proposed a dual-spaces criterion for the exact identification of critical states Liu-Critical , which overcomes the intrinsic limitations of a single-space description. Within this framework, critical states are uniquely identified by vanishing Lyapunov exponents in both position and momentum space, exhibiting a multifractal yet nonergodic form of delocalization distinct from conventional extended states. By contrast, extended and localized states necessarily exhibit an intrinsic asymmetry between the two conjugate spaces: when the Lyapunov exponent vanishes in one representation, it remains finite in the dual space, as illustrated in Fig. 1(a), (b), (e), and (f). Critical states, on the other hand, are uniquely characterized by the absence of exponential localization in both position and momentum space, as shown in Fig. 1(c) and (d).
The concept of dual-spaces invariance of critical states was first introduced by Liu and Xia in February 2023 on arXiv Liu-Critical , providing a compact and physically transparent framework for classifying eigenstates in Anderson localization. Its applicability has since been supported in a variety of systems, including flat-band lattices quasi1 and Fibonacci topological insulators quasi2 . In March 2025, work quasi3 claimed to characterize critical states through incommensurate zeros of wave functions under dual transformation, following the dual-spaces conceptual framework first introduced by Liu and Xia. These developments underscore the growing interest in dual-spaces approaches to criticality and highlight the foundational role of the dual-spaces invariance framework originally established by Liu et al., which predates and provides the conceptual basis for these subsequent studies.
A key advantage of the Liu–Xia criterion is that it allows for an exact identification of critical states, rather than relying on purely numerical multifractal analysis. Within the multifractal framework, different wave functions are characterized by distinct regimes of the scaling index $\Gamma$ You : in the thermodynamic limit, numerical results for the minimal scaling index quasi4 show that extended states satisfy $\Gamma_{\min}\to 1$, localized states obey $\Gamma_{\min}\to 0$, whereas critical states are characterized by intermediate values $0<\Gamma_{\min}<1$.
By employing the transfer-matrix method (see Appendix A) to compute the Lyapunov exponents $\gamma$ and $\gamma_{m}$ in position and momentum space, respectively, critical states can be unambiguously identified by the simultaneous vanishing of both exponents,
$\gamma(E,V)=\gamma_{m}(E,V)=0,$ (1)
where $E$ denotes the eigenenergy and $V$ the modulation strength. This compact criterion yields an exact characterization of critical states, whose multifractal nature is not readily captured by conventional single-parameter descriptions, as $\Gamma_{\min}$ does not assume a unique value but instead spans the continuous interval $(0,1)$. A unified framework for classifying Anderson localization physics is summarized in Table 1.
Table 1: Classification of typical eigenstates in disordered quantum systems based on the Liu–Xia dual-spaces criterion Liu-Critical .
States Position-Momentum Delocalization Multifractal $\Gamma_{\min}$
Extended $\gamma=0,~\gamma_{m}>0$ Position Absent 1
Localized $\gamma>0,~\gamma_{m}=0$ Momentum Absent 0
Critical $\gamma=\gamma_{m}=0$ Both Present $(0,1)$
We stress that the dual-spaces invariance discussed here refers to the invariance of localization characteristics at the level of scaling behavior and statistical properties, rather than an exact equality of critical wave functions. In systems with self-dual symmetry, the critical states at the self-dual point remain invariant under the position-momentum transformation. In contrast, without self-dual symmetry, a critical state in position space is mapped onto a distinct critical state in momentum space, while preserving the same scaling and statistical characteristics. This demonstrates that the dual-spaces invariance implied by the Liu–Xia criterion Liu-Critical is a genuinely nontrivial result. Rather than arising from self-duality that relates identical critical states, it reveals universal features shared by distinct critical states in position and momentum spaces, establishing a fundamentally new understanding of criticality in Anderson localization.
A natural question then arises: beyond Lyapunov exponents, does the dual-space characterization of critical states extend to other physical observables? In this work, we demonstrate that additional experimentally accessible quantities, most notably the inverse participation ratio, exhibit comparable scaling behavior in position and momentum space for critical states. Rather than an exact invariance, this correspondence should be understood at the level of scaling properties and statistical trends. Our results therefore complement the Liu–Xia criterion and provide a practical route for identifying multifractal critical states in both theoretical and experimental settings.
II Phenomenological Analysis
Beyond Lyapunov exponents Longhi1 ; Xia ; Jitomirskaya ; Avila , the inverse participation ratio (IPR) provide complementary information on wave-function localization IPR ; IPR1 ; IPR2 ; IPR3 . The IPR is defined as
$\mathrm{IPR}=\sum_{n=1}^{L}|\psi_{n}|^{4},$ (2)
for normalized wave functions, and quantifies the degree of spatial confinement of the probability density. Extended states exhibit $\mathrm{IPR}\sim L^{-1}$ in the thermodynamic limit, whereas exponentially localized states are characterized by a finite IPR. Closely related quantities, such as the participation entropy, have also been widely employed as probes of localization and multifractality in both theoretical and experimental studies Fan1 ; Fan2 ; Optical1 ; Optical2 ; gg0 ; gg1 ; gg2 .
For exponentially localized states, the finiteness of the IPR is consistent with the presence of a nonzero Lyapunov exponent. However, for critical states the situation is fundamentally different: their wave functions exhibit power-law decay and strong multifractal fluctuations, which preclude a description in terms of a single characteristic localization length. Consequently, no exact analytical relation between the IPR and the Lyapunov exponent can be generally established for critical states. Nevertheless, from a statistical perspective, the IPR is expected to exhibit behavior analogous to that of the Lyapunov exponent.
Motivated by the Liu–Xia criterion Liu-Critical based on the Lyapunov exponent, we now investigate how the localization properties encoded in the IPR transform between position and momentum spaces. Although the IPR is inherently basis dependent and therefore not invariant under Fourier transformation (see the Appendix B), the dual-spaces invariance underlying the Liu–Xia criterion naturally suggests that critical states should exhibit the same scaling tendency in the two representations. This expectation can be formulated in an operational form as
$\mathrm{IPR}\sim\mathrm{IPR}_{m},$ (3)
where $\mathrm{IPR}_{m}$ denotes the inverse participation ratio evaluated in momentum space, and the symbol $\sim$” signifies an asymptotic correspondence at the level of scaling behavior, rather than a strict equality. Equation (3) should therefore be interpreted as a phenomenological signature of criticality, rather than an exact invariant under Fourier transformation.
By contrast, extended and exponentially localized states generically exhibit a pronounced asymmetry between position and momentum spaces, reflected in markedly different scaling behavior of $\mathrm{IPR}$ and $\mathrm{IPR}_{m}$. Critical states, in comparison, are distinguished by a reduced dual-space asymmetry and comparable localization characteristics in both representations, consistent with their multifractal and self-similar nature. On this basis, Table 2 summarizes a phenomenological classification of eigenstates in disordered quantum systems.
Table 2: Phenomenological classification of eigenstates via the IPRs for finite-size systems.
States Position-Momentum Delocalization Multifractal $\Gamma_{\min}$
Extended $\rm IPR\sim 0,\rm IPR_{m}>0$ Position Absent 1
Localized $\rm IPR>0,\rm IPR_{m}\sim 0$ Momentum Absent 0
Critical $\rm IPR\sim\rm IPR_{m}$ Both Present $(0,1)$
III Numerical verification
To verify previous predictions, we perform numerical simulations to investigate critical states in two representative quasiperiodic models. These systems are chosen because their Hamiltonians can be represented exactly in both position and momentum spaces, making them ideal for analysis.
III.1 Aubry-André-Harper model
The Aubry-André-Harper (AAH) model AAH1 ; AAH2 represents a fundamental paradigm for studying localization phenomena in quasiperiodic systems. In position space, its Hamiltonian takes the form of a discrete Schrödinger equation:
$\psi_{n+1}+\psi_{n-1}+V\cos(2\pi\alpha n+\theta)\psi_{n}=E\psi_{n}.$ (4)
where $\psi_{n}$ is the wave function amplitude of the particle at site $n$, $V_{n}=V\cos(2\pi\alpha n+\theta)$ is the quasiperiodic potential, and $E$ is the energy eigenvalue of the particle. $\alpha$ is an irrational number, usually taken to be the golden ratio $\alpha=\frac{{\sqrt{5}-1}}{2}$, which is a typical choice for quasiperiodicity. $\theta$ is a phase factor that can be adjusted but does not affect the general results (typically set to zero for simplicity).
The AAH model exhibits a remarkable quantum phase transition driven by the quasiperiodic potential strength $V$. In the weak disorder regime ($V<2$), the system maintains extended Bloch-like eigenstates that span the entire lattice, giving rise to metallic transport properties characterized by particle delocalization. Conversely, above the critical threshold ($V>2$), the strong quasiperiodic potential induces Anderson localization, where all eigenstates become exponentially confined in space, completely suppressing particle diffusion. The transition at the critical potential strength $V=2$ represents a unique self-dual point where the system exhibits scale-invariant eigenstates with multifractal characteristics.
Utilizing the Fourier transform $\psi_{n}=\sum_{k}e^{-i2\pi\alpha nk}\phi_{k}$, the Hamiltonian of the AAH model in momentum space can be readily expressed,
$\frac{V}{2}(\phi_{k+1}+\phi_{k-1})+2\cos(2\pi\alpha k+\vartheta)\phi_{k}=E\phi_{k}.$ (5)
Thus, based on Avila’s global theory Avila1 ; Avila2 , the Lyapunov exponent in position space for the AAH model is given by $\gamma=\ln(V/2)$, while in momentum space it is $\gamma_{m}=\ln(2/V)$. Applying the Liu–Xia criterion by setting $\gamma=\gamma_{m}$, we find that the multifractal critical states occur precisely at the potential strength $V=2$. This result is in exact agreement with that obtained from the earlier self-duality relation.
III.2 Quasiperiodic-Nonlinear-Eigenproblem
The AAH model is well-known for its critical states at the self-dual point. So, does there exist a model without self-dual symmetry that still hosts critical states over a broad parameter range? Liu and Xia introduced such a model Liu-Critical , which is non-Hermitian. The construction of their model was inspired by the generalized eigenvalue problem. Mathematically, an eigenvalue problem takes the form:
$\hat{H}\left|\psi\right\rangle=E\hat{B}\left|\psi\right\rangle,$ (6)
where $\hat{H}$ denotes the Hamiltonian operator and $E$ represents the corresponding eigenvalue. The mathematical structure of Eq. (6) exhibits a fundamental dichotomy based on the form of $\hat{B}$. When $\hat{B}=I$, the equation reduces to a conventional linear eigenvalue problem. However, when $\hat{B}$ departs from the identity matrix, the problem transforms into either a nonlinear or generalized eigenvalue problem, introducing rich mathematical complexity absent in standard quantum systems.
Thus, the Quasiperiodic-Nonlinear-Eigenproblem (QNE) model Liu-Critical in momentum space can be formulated as
$\{2\cos[2\pi\alpha(k+1)]+V\}{\psi_{k+1}}+$ (7)
$\{2\cos[2\pi\alpha(k-1)]-V\}\psi_{k-1}=E(\psi_{k+1}+\psi_{k-1}).$
Non-Hermiticity in the QNE model arises through a mechanism distinct from the three well-established paradigms NH1 ; NH3 ; NH4 : (i) Non-reciprocal hopping (asymmetric tunneling amplitudes), (ii) Complex momentum (gain/loss in reciprocal hopping terms), (iii) Complex on-site potential (gain/loss localized at lattice sites). Instead, the QNE model derives its non-Hermitian character from nonlinear eigenvalue terms, representing an alternative mechanism for generating non-Hermitian spectra, distinct from the conventional paradigms based on asymmetric hopping, gain/loss terms, or complex onsite potentials. This unconventional structure leads to an enlarged critical regime: unlike the AAH model - where critical states appear only at the self-dual point $V=2$ - the QNE model sustains critical states across a broad parameter range $0<V\leq 2$.
After Fourier transformation, see the Appendix C for detailed steps, the real-space QNE Hamiltonian admits exact diagonalization as a linear eigenvalue problem, dramatically reducing computational cost:
$\phi_{n+1}+\phi_{n-1}+iV\tan(2\pi\alpha n)\phi_{n}=E\phi_{n}.$ (8)
where $i$ represents the imaginary unit, and disregarding the phase factor.
Figure 2: (Color online) Panels (a1) and (a2) depict the wave function of the AAH model (with $V=2$, $E=0.0009$) in real and momentum spaces, respectively. Panels (b1) and (b2) display the analogous results for the QNE model at $V=1$, $E=0.4$. The system size is fixed at $L=987$ lattice sites. Notably, the AAH model exhibits self-dual symmetry, under which the critical states remain invariant between position and momentum spaces. By contrast, the QNE model does not possess self-dual symmetry; as a result, a critical state in position space is mapped onto a distinct critical state in momentum space.
We emphasize that the nonlinearity manifests itself in the generalized eigenvalue structure in momentum space, while the real-space representation admits a linear non-Hermitian Hamiltonian. With the explicit forms of the Hamiltonian Eq. (7) and Eq. (8) in both real and momentum space, we can quantitatively study the characteristics of the wave functions in these two representations for the same eigenvalue.
III.3 Verification results
Figure 3: Mean inverse participation ratios in position space (MIPR) and momentum space (MIPRm) for quasiperiodic models. (a) AAH model: the two quantities exhibit markedly different finite-size scaling behaviors away from the critical point $V=2$, while they become comparable at criticality. The data are obtained from all eigenstates for system sizes L=500, 1000, 2000L=500,\;1000,\;2000. (b) QNE model: in the regime $0<V<2$, both MIPR and MIPRm show strong fluctuations and similar scaling tendencies, consistent with critical behavior over an extended parameter range. The data are obtained from six eigenstates near $E=0.00001$ for system sizes L=500, 1000, 2000L=500,\;1000,\;2000.
We perform direct numerical diagonalization for systems with $L=987$ lattice sites to obtain the eigenvalues and eigenstates of both the AAH and QNE models. The numerical results are presented in Fig. 2. As shown in panels (a1), (a2) and (b1), (b2), the critical states in both models exhibit simultaneous delocalization, multifractality, and self-similarity in both position space ($|\psi|$) and momentum space ($|\phi|$). These results confirm the theoretical assertion that the wave function of the critical state retains a certain invariance under duality transformation.
Figure 3 presents a quantitative comparison between the mean inverse participation ratios in position space (MIPR) and momentum space (MIPRm) for the two representative quasiperiodic models considered in this work. To reduce sample-to-sample fluctuations, the MIPR is obtained by averaging the IPR over a set of eigenstates near a fixed energy. For the AAH model, we take the number of sampled states to scale with the system size, whereas for the computationally more demanding QNE model we restrict the averaging to a small number of eigenstates.
As shown in Fig. 3(a), the AAH model exhibits a pronounced asymmetry between position and momentum space away from the critical point. In the regime $V<2$, the real-space MIPR displays strong system-size dependence, consistent with extended states, while the momentum-space MIPRm approaches a size-independent value, indicative of localization in momentum space. The opposite behavior is observed for $V>2$. At the self-dual point $V=2$, corresponding to criticality, the two quantities become comparable in magnitude and exhibit similar scaling behavior, in accordance with the dual-space correspondence discussed in Sec. II.
The behavior of the QNE model, shown in Fig. 3(b), is qualitatively different. For $V>2$, the eigenstates are exponentially localized in real space, leading to a weak size dependence of the MIPR, while the momentum-space MIPRm exhibits pronounced scaling with system size. In contrast, in the regime $0<V<2$, both MIPR and MIPRm display strong fluctuations and nontrivial size dependence. Rather than approaching well-defined plateaus characteristic of extended or localized phases, the two quantities exhibit comparable scaling tendencies, which is consistent with the presence of critical states over an extended parameter range.
We emphasize that, due to finite-size effects and the basis dependence of the IPR, an exact equality between MIPR and MIPRm is neither expected nor required. Instead, the numerical results demonstrate that critical states are characterized by comparable scaling behavior between position and momentum space. These observations provide direct numerical support for the dual-space correspondence formulated in Sec. II.
IV Experimental feasibility
The dual-space correspondence discussed in this work is particularly well suited for experimental verification in ultracold atomic systems, where both real-space and momentum-space observables can be accessed with high precision. One-dimensional quasiperiodic lattices realizing AAH-type Hamiltonians have been extensively implemented using optical lattices with superimposed incommensurate potentials g1 ; gg9 .
In such setups, the real-space density profile $|\psi_{n}|^{2}$ can be directly measured via in situ fluorescence or absorption imaging, enabling the extraction of the real-space inverse participation ratio $\mathrm{IPR}$. Complementarily, the momentum-space distribution $|\phi_{k}|^{2}$ can be obtained through standard time-of-flight (TOF) measurements. After suddenly switching off the trapping and lattice potentials, the atomic cloud undergoes ballistic expansion, and in the long-TOF limit the measured density profile is proportional to the initial momentum distribution. By discretizing the TOF images into momentum bins determined by the imaging resolution, the momentum-space inverse participation ratio $\mathrm{IPR}_{m}$ can be evaluated in a directly analogous manner.
Within this experimental framework, extended and localized states exhibit a pronounced asymmetry between position and momentum space, whereas critical states are expected to display comparable localization characteristics in both representations. Although finite-size effects and experimental resolution prevent an exact equality between $\mathrm{IPR}$ and $\mathrm{IPR}_{m}$, the scaling-level correspondence identified in this work should remain observable within realistic system sizes. Our findings thus establish a practical framework for diagnosing multifractal critical states with currently available cold-atom quantum simulators.
V Summary
In summary, we have investigated critical states in one-dimensional quasiperiodic systems from a dual-space perspective, focusing on the relationship between position and momentum representations. By combining analytical considerations with numerical simulations of representative models, including the AAH model and a quasiperiodic nonlinear eigenproblem, we have shown that critical states exhibit closely related localization characteristics in the two dual spaces. This behavior contrasts sharply with extended or localized states, which display strong asymmetry between real and momentum space.
Beyond Lyapunov exponents, we have demonstrated that the inverse participation ratio show comparable scaling behavior in position and momentum space for critical states. Although this correspondence does not imply an exact invariance at the level of wave-function moments, it provides a robust and practical criterion for identifying multifractal criticality in finite systems.
Our results highlight the usefulness of dual-space analysis as an operational tool for diagnosing critical states in quasiperiodic lattices. An important open question is whether similar dual-space correspondence persists in higher-dimensional or genuinely disordered systems, where the structure of momentum space is more complex. Addressing this issue would further clarify the scope and general applicability of the dual-space characterization of critical states.
Acknowledgements.
This work was supported by the Natural Science Foundation of Nanjing University of Posts and Telecommunications (Grants No. NY223109).
Appendix A Computation of the Lyapunov exponents $\gamma$ and $\gamma_{m}$ via transfer-matrix methods
1. Real-space Lyapunov exponent $\gamma$
For a given energy $E$ and a fixed realization of the quasiperiodic or disordered potential, the Lyapunov exponent $\gamma(E)$ characterizes the asymptotic exponential growth rate of wave-function amplitudes along the lattice. It can be computed efficiently using the standard transfer-matrix method. Starting from an initial vector $\mathbf{v}_{0}$ (e.g., $(1,0)^{\mathrm{T}}$), we iteratively apply the site-dependent $2\times 2$ transfer matrix $\mathbf{T}_{n}(E)$,
$\mathbf{v}_{n+1}=\mathbf{T}_{n}(E)\,\mathbf{v}_{n}.$ (9)
After $N$ iterations, the total transfer matrix reads $\boldsymbol{\Pi}_{N}(E)=\mathbf{T}_{N-1}\cdots\mathbf{T}_{1}\mathbf{T}_{0}$. The Lyapunov exponent is defined as
$\gamma(E)=\lim_{N\to\infty}\frac{1}{N}\ln\left\|\boldsymbol{\Pi}_{N}(E)\right\|,$ (10)
where $\|\cdot\|$ denotes a suitable vector norm. In numerical calculations, the vector is periodically renormalized to avoid overflow, and $\gamma(E)$ is extracted from the average logarithmic growth rate over a sufficiently long iteration ($N\sim 10^{5}$).
2. Momentum-space Lyapunov exponent $\gamma_{m}$
The momentum-space Lyapunov exponent $\gamma_{m}$ is computed using an analogous transfer-matrix approach formulated in momentum space. After Fourier transforming the real-space eigenvalue problem, one obtains a linear recurrence relation for the momentum-space amplitudes $\tilde{\psi}_{k}$. For quasiperiodic potentials with a single incommensurate frequency, this recurrence typically couples momentum modes forming an effective one-dimensional chain. Truncating the coupling beyond a sufficiently large cutoff leads to a finite-difference equation of the form
$a_{k}\tilde{\psi}_{k+Q}+b_{k}\tilde{\psi}_{k}+c_{k}\tilde{\psi}_{k-Q}=E\tilde{\psi}_{k},$ (11)
where $Q$ denotes the characteristic momentum spacing. This equation can be recast into a $2\times 2$ transfer-matrix form,
$\begin{pmatrix}\tilde{\psi}_{k+Q}\\ \tilde{\psi}_{k}\end{pmatrix}=\mathbf{M}_{k}(E)\begin{pmatrix}\tilde{\psi}_{k}\\ \tilde{\psi}_{k-Q}\end{pmatrix}$, (12)
with the matrix elements of $\mathbf{M}_{k}(E)$ determined by $a_{k}$, $b_{k}$, and $c_{k}$. Iterating this relation while periodically renormalizing the vector yields the momentum-space Lyapunov exponent,
$\gamma_{m}(E)=\lim_{K\to\infty}\frac{1}{K}\ln\left\|\boldsymbol{\Pi}^{(m)}_{K}(E)\right\|,$ (13)
where $\boldsymbol{\Pi}^{(m)}_{K}(E)$ denotes the product of momentum-space transfer matrices over $K$ steps.
In both representations, the Lyapunov exponent quantifies the exponential sensitivity of the wave-function amplitudes to lattice or momentum index. The difference lies solely in the space where the linear recurrence is defined and in the specific form of the corresponding transfer matrices.
Appendix B Remark: Basis dependence of the IPR versus Lyapunov exponents
It is instructive to clarify the distinct roles played by the Lyapunov exponent and the inverse participation ratio (IPR) in the characterization of localization properties. The Lyapunov exponent is defined through the asymptotic exponential growth (or decay) rate of solutions to a linear recurrence relation and is therefore an intrinsic dynamical quantity associated with the transfer-matrix structure of the system. As such, it does not depend on the choice of basis used to represent the wave function, but rather on the underlying recursive dynamics in a given representation (real or momentum space).
By contrast, the IPR explicitly depends on the expansion coefficients of a wave function in a chosen basis and quantifies the degree of sparsity or concentration of probability weight with respect to that representation. Since higher-order moments of wave-function amplitudes are not invariant under general unitary transformations, the IPR is inherently basis dependent and generally takes different values in position and momentum space.
This distinction is essential for the present work. While the dual-space vanishing of Lyapunov exponents provides a representation-independent criterion for criticality, the emergence of closely related scaling behavior of the IPR in dual spaces for critical states constitutes a nontrivial and physically meaningful observation. It reflects the fact that, despite the basis dependence of the IPR, critical states exhibit reduced asymmetry between position and momentum representations at the level of scaling properties.
Appendix C Fourier transformation for the QNE model
We demonstrate that the real-space equation
$\phi_{n+1}+\phi_{n-1}+iV\tan(2\pi\alpha n)\phi_{n}=E\phi_{n}$ (14)
is transformed, under a discrete Fourier transform, into the momentum-space equation
$\bigl[2\cos\bigl(2\pi\alpha(k+1)\bigr)+V\bigr]\psi_{k+1}$ (15)
$\quad+\bigl[2\cos\bigl(2\pi\alpha(k-1)\bigr)-V\bigr]\psi_{k-1}=E(\psi_{k+1}+\psi_{k-1}),$
1. Preliminaries
To eliminate the nonlocal structure of $\tan x=\sin x/\cos x$, we multiply Eq. (14) by $\cos(2\pi\alpha n)$, yielding
$\cos(2\pi\alpha n)\phi_{n+1}+\cos(2\pi\alpha n)\phi_{n-1}$ (16)
$\quad+iV\sin(2\pi\alpha n)\phi_{n}=E\cos(2\pi\alpha n)\phi_{n}.$
We define the discrete Fourier transform
$\psi_{k}=\sum_{n}e^{-i2\pi\alpha kn}\phi_{n},\qquad\phi_{n}=\sum_{k}e^{i2\pi\alpha kn}\psi_{k}.$ (17)
2. Transformation of individual terms
Using standard trigonometric identities and Fourier algebra, one finds
$\sum_{n}e^{-i2\pi\alpha kn}\cos(2\pi\alpha n)\phi_{n\pm 1}$ $=\cos\bigl[2\pi\alpha(k\pm 1)\bigr]\psi_{k\pm 1},$ (18)
$\sum_{n}e^{-i2\pi\alpha kn}\sin(2\pi\alpha n)\phi_{n}$ $=\frac{1}{2i}\bigl(\psi_{k-1}-\psi_{k+1}\bigr),$ (19)
$\sum_{n}e^{-i2\pi\alpha kn}\cos(2\pi\alpha n)\phi_{n}$ $=\frac{1}{2}\bigl(\psi_{k-1}+\psi_{k+1}\bigr).$ (20)
3. Momentum-space equation
Substituting the above relations into the Fourier transform of Eq. (16), we arrive at
$\cos\bigl[2\pi\alpha(k+1)\bigr]\psi_{k+1}+\cos\bigl[2\pi\alpha(k-1)\bigr]\psi_{k-1}$ (21)
$\quad+\frac{V}{2}(\psi_{k-1}-\psi_{k+1})=\frac{E}{2}(\psi_{k-1}+\psi_{k+1}),$
which can be rearranged into the compact form given in Eq. (15). This completes the derivation of the momentum-space representation of the QNE model.

### v6

@@
-Dual-Spaces Invariance as a Universal Criterion for Identifying Multifractal Critical States
+Dual-Space Invariance as a Universal Criterion for Multifractal Critical States
 Abstract
-In Anderson localization physics, eigenstates of disordered quantum systems are commonly classified as extended, localized, or critical, distinguished by their distinct spatial structures in real space.
-While critical states are known to exhibit multifractal characteristics, a precise and operational criterion for characterizing critical states remains an open challenge.
-In this work, we address this challenge by revisiting criticality from a dual-spaces perspective that treats position and momentum representations on equal footing.
-Building on the Liu–Xia criterion, which characterizes critical states by the simultaneous vanishing of Lyapunov exponents ($\gamma=\gamma_{m}=0$) in both spaces, we show that this dual-spaces characterization captures an essential feature of critical states that is not limited to Lyapunov exponents.
-In particular, through numerical simulations, we demonstrate that the inverse participation ratio exhibits closely related scaling behavior in position and momentum space for critical states.
-This position-momentum correspondence clearly distinguishes critical states from extended or localized ones, which instead display a pronounced asymmetry between the two representations.
-Our results establish a robust and universal framework for precisely characterizing multifractal critical states in disordered quantum systems, and provide practical guidance for their identification in current quantum simulation platforms.
+In Anderson localization, eigenstates of disordered quantum systems are broadly classified as extended, localized, or critical.
+Although critical states exhibit multifractal character, a precise and operational criterion for their identification remains an open challenge, as Lyapunov exponents in real space cannot uniquely distinguish them from extended states.
+Here we address this challenge by asserting that critical states are uniquely characterized by an emergent dual-space invariance between position and momentum space.
+Building on the Liu–Xia criterion of the simultaneous vanishing of Lyapunov exponents ($\gamma=\gamma_{m}=0$), we show that this dual-space invariance extends beyond Lyapunov exponents and governs wavefunction scaling, revealing a fundamental property inaccessible from either space alone.
+Through numerical simulations, we demonstrate that the inverse participation ratio exhibits matching scaling behavior in position and momentum space for critical states, in sharp contrast to extended and localized states, which display a pronounced asymmetry between the two spaces.
+This dual-space invariance provides a direct, robust, and universal criterion for identifying multifractal critical states.
+Our results establish a fundamental principle of Anderson criticality and open new avenues for its detection in modern quantum simulation platforms.
 I Introduction
@@
 Critical states–also referred to as multifractal states–occupy an intermediate regime between extended and localized phases You ; g1 ; zhou .
-Unlike extended states, which exhibit nearly uniform probability distributions, or localized states, which decay exponentially, critical states typically display power-law spatial decay Biddle ; Zhang ; q2 .
+Unlike extended states, which exhibit nearly uniform probability distributions, or localized states, which decay exponentially, critical states typically display power-law spatial decay Zhang ; q2 .
 Their hallmark features include multifractality, characterized by a continuous spectrum of anomalous scaling exponents q1 ; q4 ; gg10 ; gg11 , as well as scale invariance and statistical self-similarity arising from the competition between localization and delocalization gg12 ; q5 ; q6 .
-In paradigmatic settings such as the three-dimensional Anderson model, critical states emerge at mobility edges separating extended and localized spectral regions Yao1 ; Yao2 .
+In paradigmatic settings such as the three-dimensional Anderson model, critical states occur at mobility edges separating extended and localized regimes NH3 ; Yao1 ; Yao2 .
 Quasiperiodic systems cai ; Slager ; SSHQ ; gg3 ; Ayan1 ; Ayan2 provide an alternative platform for realizing critical states, even in the absence of true randomnes.
@@
 In previous theoretical descriptions Jiang1 ; Jiang2 ; gg5 ; gg6 , the asymptotic localization properties of single-particle eigenstates were typically characterized by a single real-space Lyapunov exponent $\gamma$.
+The Lyapunov exponent, as a dynamical indicator, signals the localization-delocalization transition.
 For exponentially localized states, the wave-function amplitudes decay as $|\psi_{n}|\propto e^{-\gamma|n-n_{0}|}$, where $n_{0}$ denotes the localization center and $\gamma>0$ quantifies the inverse localization length gg9 ; Duncan .
-A finite Lyapunov exponent $\gamma>0$ signals exponential localization, whereas $\gamma=0$ indicates the absence of exponential decay.
-It is important to emphasize, however, that a vanishing Lyapunov exponent $\gamma=0$ in real space does not uniquely characterize extended states: both truly extended states and critical (multifractal) states fall into the non-exponentially localized class.
+A finite Lyapunov exponent $\gamma>0$ signals exponential localization, whereas $\gamma=0$ indicates delocalization.
+It is important to emphasize, however, that a vanishing Lyapunov exponent $\gamma=0$ in real space does not uniquely characterize extended states: both truly extended states and critical (multifractal) states fall into the non-exponentially localized class and exhibit delocalization.
 Consequently, a single-space description based solely on the real-space Lyapunov exponent suffers from intrinsic limitations when applied to critical states Liu-Anomalous .
 Figure 1: (Color online) Schematic illustration of typical eigenstates in position and momentum space.
-Panels (a), (c), and (e) show extended, critical, and localized states in position space, respectively, while panels (b), (d), and (f) show their corresponding momentum-space distributions.
-In contrast to conventional approaches that emphasize real-space multifractality, a dual-space perspective highlights that critical states are distinguished by the absence of exponential localization in both representations.
-A useful conceptual perspective is offered by Bohr’s complementarity principle MQM .
-Position (real) and momentum spaces constitute a pair of Fourier-conjugate representations: while position space captures the spatial distribution of particles, momentum space encodes their reciprocal structure.
-Extended and localized states typically display strongly asymmetric localization properties in these two representations.
-This observation naturally raises the question of whether critical states, which are self-similar by construction, may exhibit a more symmetric behavior under position-momentum transformation.
-Motivated by this consideration, Liu and Xia proposed a dual-spaces criterion for the exact identification of critical states Liu-Critical , which overcomes the intrinsic limitations of a single-space description.
-Within this framework, critical states are uniquely identified by vanishing Lyapunov exponents in both position and momentum space, exhibiting a multifractal yet nonergodic form of delocalization distinct from conventional extended states.
-By contrast, extended and localized states necessarily exhibit an intrinsic asymmetry between the two conjugate spaces: when the Lyapunov exponent vanishes in one representation, it remains finite in the dual space, as illustrated in Fig.
+Panels (a), (c), and (e) show extended, critical, and localized states in position space, respectively, while panels (b), (d), and (f) display the corresponding momentum-space distributions.
+From the dual-space perspective, critical states are uniquely distinguished by the absence of exponential localization in both spaces.
+This dual absence establishes the Lyapunov exponent as a duality invariant of multifractal critical states.
+A useful conceptual perspective is provided by the Fourier uncertainty principle MQM , which dictates that localization in one representation necessarily implies delocalization in its Fourier-dual counterpart.
+Consequently, localized and extended states exhibit strongly asymmetric localization properties in position (real) and momentum spaces.
+This naturally raises the question of whether critical states, owing to their self-similar nature, may display a more symmetric behavior under position-momentum transformation.
+Motivated by this consideration, Liu and Xia proposed a dual-space criterion for the exact identification of critical states Liu-Critical , overcoming the intrinsic limitations of single-representation descriptions.
+Specifically, they proposed that critical states lack exponential localization in both position and momentum space, exhibiting vanishing Lyapunov exponents and multifractal, nonergodic delocalization.
+By contrast, extended and localized states necessarily exhibit an inherent asymmetry between the two conjugate spaces: delocalization in one representation is accompanied by localization in the dual one, as illustrated in Fig.
 1(a), (b), (e), and (f).
-Critical states, on the other hand, are uniquely characterized by the absence of exponential localization in both position and momentum space, as shown in Fig.
+Critical states, however, are uniquely characterized by invariant Lyapunov exponents, indicating delocalization in both position and momentum space, as shown in Fig.
 1(c) and (d).
-The concept of dual-spaces invariance of critical states was first introduced by Liu and Xia in February 2023 on arXiv Liu-Critical , providing a compact and physically transparent framework for classifying eigenstates in Anderson localization.
+This invariance of critical states under the Fourier transform, or dual-space transformation, is termed dual-space invariance.
+Dual-space invariance of critical states was first introduced by Liu and Xia in an arXiv preprint in February 2023 Liu-Critical .
 Its applicability has since been supported in a variety of systems, including flat-band lattices quasi1 and Fibonacci topological insulators quasi2 .
-In March 2025, work quasi3 claimed to characterize critical states through incommensurate zeros of wave functions under dual transformation, following the dual-spaces conceptual framework first introduced by Liu and Xia.
-These developments underscore the growing interest in dual-spaces approaches to criticality and highlight the foundational role of the dual-spaces invariance framework originally established by Liu et al., which predates and provides the conceptual basis for these subsequent studies.
-A key advantage of the Liu–Xia criterion is that it allows for an exact identification of critical states, rather than relying on purely numerical multifractal analysis.
-Within the multifractal framework, different wave functions are characterized by distinct regimes of the scaling index $\Gamma$ You : in the thermodynamic limit, numerical results for the minimal scaling index quasi4 show that extended states satisfy $\Gamma_{\min}\to 1$, localized states obey $\Gamma_{\min}\to 0$, whereas critical states are characterized by intermediate values $0<\Gamma_{\min}<1$.
-By employing the transfer-matrix method (see Appendix A) to compute the Lyapunov exponents $\gamma$ and $\gamma_{m}$ in position and momentum space, respectively, critical states can be unambiguously identified by the simultaneous vanishing of both exponents,
+In March 2025, Ref.
+quasi3 claimed to characterize critical states through generalized incommensurate zeros in real and dual spaces, within the dual-space conceptual framework first proposed by Liu and Xia Liu-Critical .
+These developments underscore the growing interest in dual-space approaches to criticality and further highlight the foundational role of the dual-space invariance framework introduced by Liu et al., which predates and provides the conceptual basis for these subsequent studies.
+The notion that critical states are delocalized in both position and momentum space provides an intuitive but oversimplified physical picture.
+This physical intuition is placed on a rigorous footing by the Liu–Xia criterion, which admits a precise mathematical formulation via the transfer-matrix method (see Appendix A).
+It identifies critical states through the simultaneous vanishing of the Lyapunov exponents in the two dual spaces,
 $\gamma(E,V)=\gamma_{m}(E,V)=0,$ (1)
-where $E$ denotes the eigenenergy and $V$ the modulation strength.
-This compact criterion yields an exact characterization of critical states, whose multifractal nature is not readily captured by conventional single-parameter descriptions, as $\Gamma_{\min}$ does not assume a unique value but instead spans the continuous interval $(0,1)$.
-A unified framework for classifying Anderson localization physics is summarized in Table 1.
-Table 1: Classification of typical eigenstates in disordered quantum systems based on the Liu–Xia dual-spaces criterion Liu-Critical .
+where $\gamma$ and $\gamma_{m}$ correspond to position and momentum space, respectively, $E$ is the eigenenergy, and $V$ the modulation strength.
+This implies that the Lyapunov exponent serves as a duality invariant of critical states, and its advantages become evident in comparison with finite-size multifractal analysis.
+According to multifractal theory You ; quasi4 , in the thermodynamic limit, the minimal scaling index $\Gamma_{\min}$ approaches $1$ for extended states and 0 for localized states, while critical states satisfy $0<\Gamma_{\min}<1$.
+However, finite-size multifractal analysis suffers from intrinsic ambiguity, as critical states with small scaling exponents are difficult to distinguish from localized states.
+The Liu–Xia criterion overcomes this fundamental limitation by providing an exact and representation-independent identification of critical states.
+This is essential because the multifractal nature of critical states cannot be uniquely characterized by a single scaling exponent, as $\Gamma_{\min}$ spans a continuous interval $(0,1)$ rather than converging to a definite value.
+Finally, a unified framework for Anderson localization is summarized in Table 1.
+Table 1: Classification of typical eigenstates in disordered quantum systems based on the Liu–Xia dual-space criterion Liu-Critical .
 States Position-Momentum Delocalization Multifractal $\Gamma_{\min}$
@@
 Critical $\gamma=\gamma_{m}=0$ Both Present $(0,1)$
-We stress that the dual-spaces invariance discussed here refers to the invariance of localization characteristics at the level of scaling behavior and statistical properties, rather than an exact equality of critical wave functions.
-In systems with self-dual symmetry, the critical states at the self-dual point remain invariant under the position-momentum transformation.
-In contrast, without self-dual symmetry, a critical state in position space is mapped onto a distinct critical state in momentum space, while preserving the same scaling and statistical characteristics.
-This demonstrates that the dual-spaces invariance implied by the Liu–Xia criterion Liu-Critical is a genuinely nontrivial result.
-Rather than arising from self-duality that relates identical critical states, it reveals universal features shared by distinct critical states in position and momentum spaces, establishing a fundamentally new understanding of criticality in Anderson localization.
-A natural question then arises: beyond Lyapunov exponents, does the dual-space characterization of critical states extend to other physical observables?
-In this work, we demonstrate that additional experimentally accessible quantities, most notably the inverse participation ratio, exhibit comparable scaling behavior in position and momentum space for critical states.
-Rather than an exact invariance, this correspondence should be understood at the level of scaling properties and statistical trends.
-Our results therefore complement the Liu–Xia criterion and provide a practical route for identifying multifractal critical states in both theoretical and experimental settings.
+A natural question then arises: beyond Lyapunov exponents, does the dual-space characterization of critical states extend to other observables?
+Here, we show that experimentally accessible quantities, particularly the inverse participation ratio, exhibit analogous scaling in position and momentum space at criticality.
+This correspondence arises at the level of scaling and statistical properties rather than exact invariance.
+Our results therefore complement the Liu–Xia criterion and provide a practical route to identifying multifractal critical states in both theoretical and experimental settings.
 II Phenomenological Analysis
-Beyond Lyapunov exponents Longhi1 ; Xia ; Jitomirskaya ; Avila , the inverse participation ratio (IPR) provide complementary information on wave-function localization IPR ; IPR1 ; IPR2 ; IPR3 .
-The IPR is defined as
+Compared with the Lyapunov exponent, the inverse participation ratio (IPR) provides complementary information on wave-function localization Longhi1 ; Xia ; Jitomirskaya ; Avila ; IPR ; IPR1 ; IPR2 ; IPR3 .
+It is defined for normalized wave functions as
 $\mathrm{IPR}=\sum_{n=1}^{L}|\psi_{n}|^{4},$ (2)
-for normalized wave functions, and quantifies the degree of spatial confinement of the probability density.
+and quantifies the spatial confinement of the probability density.
 Extended states exhibit $\mathrm{IPR}\sim L^{-1}$ in the thermodynamic limit, whereas exponentially localized states are characterized by a finite IPR.
-Closely related quantities, such as the participation entropy, have also been widely employed as probes of localization and multifractality in both theoretical and experimental studies Fan1 ; Fan2 ; Optical1 ; Optical2 ; gg0 ; gg1 ; gg2 .
-For exponentially localized states, the finiteness of the IPR is consistent with the presence of a nonzero Lyapunov exponent.
-However, for critical states the situation is fundamentally different: their wave functions exhibit power-law decay and strong multifractal fluctuations, which preclude a description in terms of a single characteristic localization length.
-Consequently, no exact analytical relation between the IPR and the Lyapunov exponent can be generally established for critical states.
-Nevertheless, from a statistical perspective, the IPR is expected to exhibit behavior analogous to that of the Lyapunov exponent.
+Related quantities, such as the participation entropy, have also been widely used to probe localization and multifractality in theoretical and experimental studies Fan1 ; Fan2 ; Optical1 ; Optical2 ; gg0 ; gg1 ; gg2 .
+For exponentially localized states, the finite IPR is consistent with a nonzero Lyapunov exponent.
+In contrast, critical states exhibit power-law decay and strong multifractal fluctuations, which preclude a description in terms of a single localization length.
+As a result, no exact analytical relation between the IPR and the Lyapunov exponent generally exists for critical states.
+Nevertheless, from a statistical perspective, the IPR is expected to display behavior analogous to that of the Lyapunov exponent.
 Motivated by the Liu–Xia criterion Liu-Critical based on the Lyapunov exponent, we now investigate how the localization properties encoded in the IPR transform between position and momentum spaces.
-Although the IPR is inherently basis dependent and therefore not invariant under Fourier transformation (see the Appendix B), the dual-spaces invariance underlying the Liu–Xia criterion naturally suggests that critical states should exhibit the same scaling tendency in the two representations.
+Although the IPR is inherently basis dependent and therefore not invariant under Fourier transformation (see the Appendix B), the dual-space invariance underlying the Liu–Xia criterion naturally suggests that critical states should exhibit the same scaling tendency in the two representations.
 This expectation can be formulated in an operational form as
@@
 Equation (3) should therefore be interpreted as a phenomenological signature of criticality, rather than an exact invariant under Fourier transformation.
-By contrast, extended and exponentially localized states generically exhibit a pronounced asymmetry between position and momentum spaces, reflected in markedly different scaling behavior of $\mathrm{IPR}$ and $\mathrm{IPR}_{m}$.
-Critical states, in comparison, are distinguished by a reduced dual-space asymmetry and comparable localization characteristics in both representations, consistent with their multifractal and self-similar nature.
+By contrast, extended and exponentially localized states typically exhibit a pronounced asymmetry between position and momentum spaces, reflected in the markedly different scaling behavior of $\mathrm{IPR}$ and $\mathrm{IPR}_{m}$.
+In comparison, critical states remain protected by dual-space symmetry and exhibit comparable localization characteristics in both representations, consistent with their multifractal and self-similar nature.
 On this basis, Table 2 summarizes a phenomenological classification of eigenstates in disordered quantum systems.
@@
 Critical $\rm IPR\sim\rm IPR_{m}$ Both Present $(0,1)$
+This demonstrates that the Lyapunov exponent itself is an exact duality invariant for multifractal critical states, whereas the IPR, owing to its intrinsic basis dependence, is not strictly duality-invariant.
+It should be noted that the dual-space invariance discussed here refers to the invariance of localization characteristics at the level of scaling behavior and statistical properties, rather than an exact equality of critical wave functions.
+In systems with self-dual symmetry, the critical states at the self-dual point remain invariant under the position-momentum transformation.
+In contrast, without self-dual symmetry, a critical state in position space is mapped onto a distinct critical state in momentum space, while preserving the same scaling and statistical characteristics.
+This demonstrates that the dual-space invariance implied by the Liu–Xia criterion Liu-Critical is a genuinely nontrivial result.
+Rather than arising from self-duality that relates identical critical states, it reveals universal features shared by distinct critical states in position and momentum spaces, establishing a fundamentally new understanding of criticality in Anderson localization.
 III Numerical verification
@@
 $\psi_{n+1}+\psi_{n-1}+V\cos(2\pi\alpha n+\theta)\psi_{n}=E\psi_{n}.$ (4)
-where $\psi_{n}$ is the wave function amplitude of the particle at site $n$, $V_{n}=V\cos(2\pi\alpha n+\theta)$ is the quasiperiodic potential, and $E$ is the energy eigenvalue of the particle.
-$\alpha$ is an irrational number, usually taken to be the golden ratio $\alpha=\frac{{\sqrt{5}-1}}{2}$, which is a typical choice for quasiperiodicity.
-$\theta$ is a phase factor that can be adjusted but does not affect the general results (typically set to zero for simplicity).
-The AAH model exhibits a remarkable quantum phase transition driven by the quasiperiodic potential strength $V$.
-In the weak disorder regime ($V<2$), the system maintains extended Bloch-like eigenstates that span the entire lattice, giving rise to metallic transport properties characterized by particle delocalization.
-Conversely, above the critical threshold ($V>2$), the strong quasiperiodic potential induces Anderson localization, where all eigenstates become exponentially confined in space, completely suppressing particle diffusion.
-The transition at the critical potential strength $V=2$ represents a unique self-dual point where the system exhibits scale-invariant eigenstates with multifractal characteristics.
+where $\psi_{n}$ is the wave function amplitude of the particle at site $n$, $V_{n}=V\cos(2\pi\alpha n+\theta)$ is the quasiperiodic potential, and $E$ is the eigenvalue.
+$\alpha$ is an irrational number, usually taken to be the golden ratio $\alpha=\frac{{\sqrt{5}-1}}{2}$.
+$\theta$ is a phase factor, typically taken to be zero.
 Utilizing the Fourier transform $\psi_{n}=\sum_{k}e^{-i2\pi\alpha nk}\phi_{k}$, the Hamiltonian of the AAH model in momentum space can be readily expressed,
 $\frac{V}{2}(\phi_{k+1}+\phi_{k-1})+2\cos(2\pi\alpha k+\vartheta)\phi_{k}=E\phi_{k}.$ (5)
-Thus, based on Avila’s global theory Avila1 ; Avila2 , the Lyapunov exponent in position space for the AAH model is given by $\gamma=\ln(V/2)$, while in momentum space it is $\gamma_{m}=\ln(2/V)$.
-Applying the Liu–Xia criterion by setting $\gamma=\gamma_{m}$, we find that the multifractal critical states occur precisely at the potential strength $V=2$.
-This result is in exact agreement with that obtained from the earlier self-duality relation.
+The AAH model exhibits a quantum phase transition driven by the quasiperiodic potential strength $V$.
+For weak quasiperiodic potential ($V<2$), all eigenstates are extended Bloch-like states, resulting in metallic transport.
+In contrast, for $V>2$, the system undergoes Anderson localization, where all eigenstates become exponentially localized, suppressing particle diffusion.
+At the critical point $V=2$, the Hamiltonians in real space [Eq.
+(4)] and momentum space [Eq.
+(5)] take identical forms.
+Consequently, the system reaches the self-dual point and exhibits multifractal eigenstates.
+The procedure for identifying multifractal critical states according to the Liu–Xia criterion is outlined as follows.
+From Avila’s global theory Avila1 ; Avila2 , the Lyapunov exponents in position and momentum spaces are given by $\gamma=\ln(V/2)$ and $\gamma_{m}=\ln(2/V)$, respectively.
+Imposing the Liu–Xia criterion, $\gamma=\gamma_{m}$, rigorously and uniquely determines the critical point at $V=2$, thereby establishing the multifractal nature of the eigenstates.
+This result is in exact agreement with that obtained from the self-duality relation, further confirming the validity and consistency of the criterion.
 III.2 Quasiperiodic-Nonlinear-Eigenproblem
@@
 $\{2\cos[2\pi\alpha(k-1)]-V\}\psi_{k-1}=E(\psi_{k+1}+\psi_{k-1}).$
-Non-Hermiticity in the QNE model arises through a mechanism distinct from the three well-established paradigms NH1 ; NH3 ; NH4 : (i) Non-reciprocal hopping (asymmetric tunneling amplitudes), (ii) Complex momentum (gain/loss in reciprocal hopping terms), (iii) Complex on-site potential (gain/loss localized at lattice sites).
-Instead, the QNE model derives its non-Hermitian character from nonlinear eigenvalue terms, representing an alternative mechanism for generating non-Hermitian spectra, distinct from the conventional paradigms based on asymmetric hopping, gain/loss terms, or complex onsite potentials.
-This unconventional structure leads to an enlarged critical regime: unlike the AAH model - where critical states appear only at the self-dual point $V=2$ - the QNE model sustains critical states across a broad parameter range $0<V\leq 2$.
-After Fourier transformation, see the Appendix C for detailed steps, the real-space QNE Hamiltonian admits exact diagonalization as a linear eigenvalue problem, dramatically reducing computational cost:
+Non-Hermiticity in the QNE model originates from nonlinear eigenvalue terms, fundamentally distinct from the three established paradigms of asymmetric hopping, gain/loss, and complex potentials NH1 ; NH4 .
+This unconventional mechanism provides a new route to non-Hermitian spectra.
+As a direct consequence, the QNE model exhibits an expanded critical regime.
+Compared with the AAH model, where critical states exist only at the self-dual point $V=2$, the QNE model supports critical states over a broad parameter range $0<V\leq 2$.
+After Fourier transformation, the real-space QNE Hamiltonian reduces to a linear eigenvalue problem that can be exactly diagonalized, dramatically reducing the computational cost (see Appendix C for details),
 $\phi_{n+1}+\phi_{n-1}+iV\tan(2\pi\alpha n)\phi_{n}=E\phi_{n}.$ (8)
 where $i$ represents the imaginary unit, and disregarding the phase factor.
-Figure 2: (Color online) Panels (a1) and (a2) depict the wave function of the AAH model (with $V=2$, $E=0.0009$) in real and momentum spaces, respectively.
-Panels (b1) and (b2) display the analogous results for the QNE model at $V=1$, $E=0.4$.
-The system size is fixed at $L=987$ lattice sites.
-Notably, the AAH model exhibits self-dual symmetry, under which the critical states remain invariant between position and momentum spaces.
-By contrast, the QNE model does not possess self-dual symmetry; as a result, a critical state in position space is mapped onto a distinct critical state in momentum space.
-We emphasize that the nonlinearity manifests itself in the generalized eigenvalue structure in momentum space, while the real-space representation admits a linear non-Hermitian Hamiltonian.
-With the explicit forms of the Hamiltonian Eq.
+Figure 2: (Color online) Panels (a1) and (a2) show the wave function of the AAH model ($V=2$, $E=0.0009$) in real and momentum space, respectively, while (b1) and (b2) show the corresponding results for the QNE model ($V=1$, $E=0.4$).
+The system size is $L=987$.
+Due to self-duality, critical states of the AAH model are invariant between real and momentum space.
+In contrast, the QNE model lacks self-duality: a critical state transforms into a distinct critical state under Fourier transformation, while their Lyapunov exponents remain identical.
+Clearly, unlike the AAH model, the QNE model lacks self-duality symmetry, and its critical states generally exhibit different structures in real and momentum spaces.
+Consequently, the self-duality relation completely fails to identify the critical states.
+In sharp contrast, the Liu–Xia criterion remains fully operative.
+Specifically, the Lyapunov exponents in real and momentum spaces can be rigorously obtained from Eq.
 (8) and Eq.
-(8) in both real and momentum space, we can quantitatively study the characteristics of the wave functions in these two representations for the same eigenvalue.
+(7), respectively, and the critical states are uniquely determined by enforcing their equality, as demonstrated in Ref.
+Liu-Critical .
+Despite the absence of any self-duality symmetry, the critical states are still dictated by the invariance of the Lyapunov exponents in dual spaces, thereby revealing the fundamental origin, conceptual advance, and universal validity of the Liu–Xia criterion.
 III.3 Verification results
-Figure 3: Mean inverse participation ratios in position space (MIPR) and momentum space (MIPRm) for quasiperiodic models.
-(a) AAH model: the two quantities exhibit markedly different finite-size scaling behaviors away from the critical point $V=2$, while they become comparable at criticality.
-The data are obtained from all eigenstates for system sizes L=500, 1000, 2000L=500,\;1000,\;2000.
-(b) QNE model: in the regime $0<V<2$, both MIPR and MIPRm show strong fluctuations and similar scaling tendencies, consistent with critical behavior over an extended parameter range.
-The data are obtained from six eigenstates near $E=0.00001$ for system sizes L=500, 1000, 2000L=500,\;1000,\;2000.
-We perform direct numerical diagonalization for systems with $L=987$ lattice sites to obtain the eigenvalues and eigenstates of both the AAH and QNE models.
-The numerical results are presented in Fig.
+Figure 3: (a) AAH model: MIPR and MIPRm exhibit distinct scaling away from the critical point $V=2$, but converge to comparable values at criticality.
+Results are averaged over all eigenstates for L=500, 1000, 2000L=500,\;1000,\;2000.
+(b) QNE model: in the regime $0<V<2$, both MIPR and MIPRm display strong fluctuations yet remain comparable, supporting criticality over a broadened parameter range.
+Results are obtained from six eigenstates near $E=0.00001$ for L=500, 1000, 2000L=500,\;1000,\;2000.
+We perform direct numerical diagonalization for systems with $L=987$ to obtain the eigenvalues and eigenstates of both the AAH and QNE models.
+The results are shown in Fig.
 2.
-As shown in panels (a1), (a2) and (b1), (b2), the critical states in both models exhibit simultaneous delocalization, multifractality, and self-similarity in both position space ($|\psi|$) and momentum space ($|\phi|$).
-These results confirm the theoretical assertion that the wave function of the critical state retains a certain invariance under duality transformation.
-Figure 3 presents a quantitative comparison between the mean inverse participation ratios in position space (MIPR) and momentum space (MIPRm) for the two representative quasiperiodic models considered in this work.
-To reduce sample-to-sample fluctuations, the MIPR is obtained by averaging the IPR over a set of eigenstates near a fixed energy.
-For the AAH model, we take the number of sampled states to scale with the system size, whereas for the computationally more demanding QNE model we restrict the averaging to a small number of eigenstates.
+As seen in panels (a1), (a2) and (b1), (b2), the critical states exhibit simultaneous delocalization, multifractality, and self-similarity in both position space ($|\psi|$) and momentum space ($|\phi|$).
+These results provide direct numerical evidence that critical wave functions obey dual-space invariance under Fourier transformation.
+Figure 3 compares the mean inverse participation ratios in position space (MIPR) and momentum space (MIPRm) for the two quasiperiodic models.
+The MIPR is computed by averaging the IPR over eigenstates within a narrow energy window to suppress sample fluctuations.
+For the AAH model, the number of sampled states scales with system size, whereas for the QNE model, the averaging is limited to a few eigenstates due to computational cost.
 As shown in Fig.
-3(a), the AAH model exhibits a pronounced asymmetry between position and momentum space away from the critical point.
-In the regime $V<2$, the real-space MIPR displays strong system-size dependence, consistent with extended states, while the momentum-space MIPRm approaches a size-independent value, indicative of localization in momentum space.
-The opposite behavior is observed for $V>2$.
-At the self-dual point $V=2$, corresponding to criticality, the two quantities become comparable in magnitude and exhibit similar scaling behavior, in accordance with the dual-space correspondence discussed in Sec.
-II.
+3(a), the AAH model exhibits strong asymmetry between position and momentum space away from criticality.
+For $V<2$, the real-space MIPR decreases with system size, consistent with extended states, while the momentum-space MIPRm remains finite, indicating localization in momentum space.
+The opposite behavior occurs for $V>2$.
+At the self-dual point $V=2$, the two quantities become comparable and follow similar scaling, reflecting dual-space invariance at criticality.
 The behavior of the QNE model, shown in Fig.
 3(b), is qualitatively different.
-For $V>2$, the eigenstates are exponentially localized in real space, leading to a weak size dependence of the MIPR, while the momentum-space MIPRm exhibits pronounced scaling with system size.
-In contrast, in the regime $0<V<2$, both MIPR and MIPRm display strong fluctuations and nontrivial size dependence.
-Rather than approaching well-defined plateaus characteristic of extended or localized phases, the two quantities exhibit comparable scaling tendencies, which is consistent with the presence of critical states over an extended parameter range.
-We emphasize that, due to finite-size effects and the basis dependence of the IPR, an exact equality between MIPR and MIPRm is neither expected nor required.
-Instead, the numerical results demonstrate that critical states are characterized by comparable scaling behavior between position and momentum space.
-These observations provide direct numerical support for the dual-space correspondence formulated in Sec.
-II.
+For $V>2$, the eigenstates are exponentially localized in real space, leading to a weak size dependence of the MIPR, while the momentum-space MIPRm exhibits pronounced system-size scaling.
+In contrast, for $0<V<2$, both MIPR and MIPRm show strong fluctuations and nontrivial size dependence.
+Rather than approaching plateaus characteristic of extended or localized phases, the two quantities exhibit comparable scaling, consistent with a critical regime.
+We emphasize that, due to finite-size effects and the basis dependence of the IPR, exact equality between MIPR and MIPRm is neither expected nor required.
+Instead, the results show that critical states are characterized by comparable scaling in position and momentum space.
+These findings provide direct numerical support for the dual-space correspondence.
 IV Experimental feasibility
-The dual-space correspondence discussed in this work is particularly well suited for experimental verification in ultracold atomic systems, where both real-space and momentum-space observables can be accessed with high precision.
-One-dimensional quasiperiodic lattices realizing AAH-type Hamiltonians have been extensively implemented using optical lattices with superimposed incommensurate potentials g1 ; gg9 .
-In such setups, the real-space density profile $|\psi_{n}|^{2}$ can be directly measured via in situ fluorescence or absorption imaging, enabling the extraction of the real-space inverse participation ratio $\mathrm{IPR}$.
-Complementarily, the momentum-space distribution $|\phi_{k}|^{2}$ can be obtained through standard time-of-flight (TOF) measurements.
-After suddenly switching off the trapping and lattice potentials, the atomic cloud undergoes ballistic expansion, and in the long-TOF limit the measured density profile is proportional to the initial momentum distribution.
-By discretizing the TOF images into momentum bins determined by the imaging resolution, the momentum-space inverse participation ratio $\mathrm{IPR}_{m}$ can be evaluated in a directly analogous manner.
-Within this experimental framework, extended and localized states exhibit a pronounced asymmetry between position and momentum space, whereas critical states are expected to display comparable localization characteristics in both representations.
-Although finite-size effects and experimental resolution prevent an exact equality between $\mathrm{IPR}$ and $\mathrm{IPR}_{m}$, the scaling-level correspondence identified in this work should remain observable within realistic system sizes.
-Our findings thus establish a practical framework for diagnosing multifractal critical states with currently available cold-atom quantum simulators.
-V Summary
-In summary, we have investigated critical states in one-dimensional quasiperiodic systems from a dual-space perspective, focusing on the relationship between position and momentum representations.
-By combining analytical considerations with numerical simulations of representative models, including the AAH model and a quasiperiodic nonlinear eigenproblem, we have shown that critical states exhibit closely related localization characteristics in the two dual spaces.
-This behavior contrasts sharply with extended or localized states, which display strong asymmetry between real and momentum space.
-Beyond Lyapunov exponents, we have demonstrated that the inverse participation ratio show comparable scaling behavior in position and momentum space for critical states.
-Although this correspondence does not imply an exact invariance at the level of wave-function moments, it provides a robust and practical criterion for identifying multifractal criticality in finite systems.
-Our results highlight the usefulness of dual-space analysis as an operational tool for diagnosing critical states in quasiperiodic lattices.
-An important open question is whether similar dual-space correspondence persists in higher-dimensional or genuinely disordered systems, where the structure of momentum space is more complex.
-Addressing this issue would further clarify the scope and general applicability of the dual-space characterization of critical states.
+The dual-space correspondence identified in this work is well suited for experimental verification in ultracold atomic systems, where both position- and momentum-space observables are directly accessible.
+One-dimensional quasiperiodic lattices realizing AAH-type Hamiltonians can be implemented using optical lattices with superimposed incommensurate potentials.
+In such setups, the real-space density distribution $|\psi_{n}|^{2}$ can be measured using in situ fluorescence or absorption imaging gg9 , allowing direct evaluation of the real-space IPR.
+The momentum-space distribution $|\phi_{k}|^{2}$ can be obtained via standard time-of-flight (TOF) measurements Biddle .
+After switching off the trapping and lattice potentials, the atomic cloud undergoes ballistic expansion, and in the long-TOF limit the measured density profile reflects the initial momentum distribution.
+By discretizing the TOF images according to the imaging resolution, the momentum-space inverse participation ratio $\mathrm{IPR}_{m}$ can be extracted in a directly analogous manner.
+Within this framework, extended and localized states exhibit pronounced asymmetry between position and momentum space, whereas critical states show comparable localization properties in both representations.
+Although finite system sizes and experimental resolution prevent exact quantitative agreement, the scaling-level correspondence predicted here remains observable under realistic conditions.
+These results demonstrate that current cold-atom experiments provide a feasible and practical route for identifying multifractal critical states through dual-space diagnostics.
+V Summary and outlook
+In summary, we have investigated critical states in one-dimensional quasiperiodic systems from a dual-space perspective.
+Using analytical arguments and numerical simulations of representative models, including the AAH model and a quasiperiodic nonlinear eigenproblem, we have shown that critical states are uniquely characterized by comparable scaling in dual spaces.
+This property sharply distinguishes them from extended and localized states, which exhibit a pronounced asymmetry between position and momentum spaces.
+We have further demonstrated that the inverse participation ratio provides a simple and practical diagnostic of this behavior: critical states exhibit comparable scaling of the inverse participation ratio in position and momentum space even in finite systems.
+This result establishes dual-space correspondence as an operational criterion for identifying multifractal criticality beyond conventional approaches based solely on a single space.
+An important open direction is to examine interacting many-body systems, where quasiperiodicity and interactions can produce many-body localized and critical phases.
+It will be particularly interesting to determine whether a similar scaling correspondence emerges between real-space and Fock-space or momentum-space structures of many-body eigenstates, and whether the Liu–Xia criterion can be extended to provide a precise characterization of many-body criticality.
 Acknowledgements.
@@
 1. Real-space Lyapunov exponent $\gamma$
-For a given energy $E$ and a fixed realization of the quasiperiodic or disordered potential, the Lyapunov exponent $\gamma(E)$ characterizes the asymptotic exponential growth rate of wave-function amplitudes along the lattice.
-It can be computed efficiently using the standard transfer-matrix method.
-Starting from an initial vector $\mathbf{v}_{0}$ (e.g., $(1,0)^{\mathrm{T}}$), we iteratively apply the site-dependent $2\times 2$ transfer matrix $\mathbf{T}_{n}(E)$,
+For a given energy $E$, the real-space Lyapunov exponent $\gamma(E)$ characterizes the asymptotic exponential growth of the wave-function amplitude along the lattice.
+It is computed using the standard transfer-matrix method.
+Starting from an initial vector $\mathbf{v}_{0}$ (e.g., $(1,0)^{\mathrm{T}}$), one iterates the site-dependent transfer matrix $\mathbf{T}_{n}(E)$,
 $\mathbf{v}_{n+1}=\mathbf{T}_{n}(E)\mathbf{v}_{n}.$ (9)
-After $N$ iterations, the total transfer matrix reads $\boldsymbol{\Pi}_{N}(E)=\mathbf{T}_{N-1}\cdots\mathbf{T}_{1}\mathbf{T}_{0}$. The Lyapunov exponent is defined as
-$\gamma(E)=\lim_{N\to\infty}\frac{1}{N}\ln\left\|\boldsymbol{\Pi}_{N}(E)\right\|,$ (10)
-where $\|\cdot\|$ denotes a suitable vector norm.
-In numerical calculations, the vector is periodically renormalized to avoid overflow, and $\gamma(E)$ is extracted from the average logarithmic growth rate over a sufficiently long iteration ($N\sim 10^{5}$).
+After $N$ steps, the total transfer matrix is $\boldsymbol{\Pi}_{N}(E)=\mathbf{T}_{N-1}\cdots\mathbf{T}_{0}$, and the Lyapunov exponent is defined as
+$\gamma(E)=\lim_{N\to\infty}\frac{1}{N}\ln\left\|\boldsymbol{\Pi}_{N}(E)\right\|.$ (10)
+In numerical implementation, periodic renormalization is applied to prevent overflow, and $\gamma(E)$ is obtained from the averaged logarithmic growth rate for large $N$ (typically $N\sim 10^{5}$).
 2. Momentum-space Lyapunov exponent $\gamma_{m}$
-The momentum-space Lyapunov exponent $\gamma_{m}$ is computed using an analogous transfer-matrix approach formulated in momentum space.
-After Fourier transforming the real-space eigenvalue problem, one obtains a linear recurrence relation for the momentum-space amplitudes $\tilde{\psi}_{k}$.
-For quasiperiodic potentials with a single incommensurate frequency, this recurrence typically couples momentum modes forming an effective one-dimensional chain.
-Truncating the coupling beyond a sufficiently large cutoff leads to a finite-difference equation of the form
+The momentum-space Lyapunov exponent $\gamma_{m}(E)$ is computed using an analogous transfer-matrix formulation in momentum space.
+After Fourier transforming the eigenvalue equation, the momentum-space amplitudes $\tilde{\psi}_{k}$ satisfy a linear recurrence relation.
+For quasiperiodic potentials with a single incommensurate frequency, this relation forms an effective one-dimensional chain,
 $a_{k}\tilde{\psi}_{k+Q}+b_{k}\tilde{\psi}_{k}+c_{k}\tilde{\psi}_{k-Q}=E\tilde{\psi}_{k},$ (11)
-where $Q$ denotes the characteristic momentum spacing. This equation can be recast into a $2\times 2$ transfer-matrix form,
-$\begin{pmatrix}\tilde{\psi}_{k+Q}\\ \tilde{\psi}_{k}\end{pmatrix}=\mathbf{M}_{k}(E)\begin{pmatrix}\tilde{\psi}_{k}\\ \tilde{\psi}_{k-Q}\end{pmatrix}$, (12)
-with the matrix elements of $\mathbf{M}_{k}(E)$ determined by $a_{k}$, $b_{k}$, and $c_{k}$.
-Iterating this relation while periodically renormalizing the vector yields the momentum-space Lyapunov exponent,
+which can be rewritten as
+$\begin{pmatrix}\tilde{\psi}_{k+Q}\\ \tilde{\psi}_{k}\end{pmatrix}=\mathbf{M}_{k}(E)\begin{pmatrix}\tilde{\psi}_{k}\\ \tilde{\psi}_{k-Q}\end{pmatrix}$. (12)
+Iterating the corresponding transfer matrices yields
 $\gamma_{m}(E)=\lim_{K\to\infty}\frac{1}{K}\ln\left\|\boldsymbol{\Pi}^{(m)}_{K}(E)\right\|,$ (13)
-where $\boldsymbol{\Pi}^{(m)}_{K}(E)$ denotes the product of momentum-space transfer matrices over $K$ steps.
-In both representations, the Lyapunov exponent quantifies the exponential sensitivity of the wave-function amplitudes to lattice or momentum index.
-The difference lies solely in the space where the linear recurrence is defined and in the specific form of the corresponding transfer matrices.
-Appendix B Remark: Basis dependence of the IPR versus Lyapunov exponents
-It is instructive to clarify the distinct roles played by the Lyapunov exponent and the inverse participation ratio (IPR) in the characterization of localization properties.
-The Lyapunov exponent is defined through the asymptotic exponential growth (or decay) rate of solutions to a linear recurrence relation and is therefore an intrinsic dynamical quantity associated with the transfer-matrix structure of the system.
-As such, it does not depend on the choice of basis used to represent the wave function, but rather on the underlying recursive dynamics in a given representation (real or momentum space).
-By contrast, the IPR explicitly depends on the expansion coefficients of a wave function in a chosen basis and quantifies the degree of sparsity or concentration of probability weight with respect to that representation.
-Since higher-order moments of wave-function amplitudes are not invariant under general unitary transformations, the IPR is inherently basis dependent and generally takes different values in position and momentum space.
-This distinction is essential for the present work.
-While the dual-space vanishing of Lyapunov exponents provides a representation-independent criterion for criticality, the emergence of closely related scaling behavior of the IPR in dual spaces for critical states constitutes a nontrivial and physically meaningful observation.
-It reflects the fact that, despite the basis dependence of the IPR, critical states exhibit reduced asymmetry between position and momentum representations at the level of scaling properties.
+where $\boldsymbol{\Pi}^{(m)}_{K}(E)$ denotes the product of momentum-space transfer matrices.
+In both cases, the Lyapunov exponent quantifies the exponential growth rate of the transfer-matrix norm, with the only difference being whether the recurrence is defined in position or momentum space.
+Appendix B Contrast Between Basis-Dependent IPR and Basis-Invariant Lyapunov Exponents
+It is instructive to contrast the distinct roles of the Lyapunov exponent and the inverse participation ratio (IPR) in characterizing localization.
+The Lyapunov exponent is defined by the asymptotic exponential growth or decay of solutions to a linear recurrence relation and thus represents an intrinsic dynamical quantity associated with the transfer-matrix structure.
+Consequently, it is independent of the choice of basis and depends only on the underlying recursive dynamics in a given representation (real or momentum space).
+In contrast, the IPR depends explicitly on the expansion coefficients of the wave function in a chosen basis and quantifies the degree of localization within that representation.
+Because higher-order moments of the wave-function amplitudes are not invariant under unitary transformations, the IPR is inherently basis dependent and generally differs between position and momentum space.
+This distinction is central to the present work.
+While the simultaneous vanishing of Lyapunov exponents in dual spaces provides a basis-invariant criterion for criticality, the emergence of comparable scaling of the IPR in dual spaces constitutes a nontrivial physical signature.
+It indicates that, despite its basis dependence, the IPR reflects the reduced asymmetry between position and momentum space at criticality.
 Appendix C Fourier transformation for the QNE model
-We demonstrate that the real-space equation
+We show that the real-space equation
 $\phi_{n+1}+\phi_{n-1}+iV\tan(2\pi\alpha n)\phi_{n}=E\phi_{n}$ (14)
-is transformed, under a discrete Fourier transform, into the momentum-space equation
+transforms into the momentum-space equation
 $\bigl[2\cos\bigl(2\pi\alpha(k+1)\bigr)+V\bigr]\psi_{k+1}$ (15)
 $\quad+\bigl[2\cos\bigl(2\pi\alpha(k-1)\bigr)-V\bigr]\psi_{k-1}=E(\psi_{k+1}+\psi_{k-1}),$
-1. Preliminaries
-To eliminate the nonlocal structure of $\tan x=\sin x/\cos x$, we multiply Eq. (14) by $\cos(2\pi\alpha n)$, yielding
+under a discrete Fourier transform.
+1. Fourier transform
+Multiplying Eq. (14) by $\cos(2\pi\alpha n)$ gives
 $\cos(2\pi\alpha n)\phi_{n+1}+\cos(2\pi\alpha n)\phi_{n-1}$ (16)
@@
 $\psi_{k}=\sum_{n}e^{-i2\pi\alpha kn}\phi_{n},\qquad\phi_{n}=\sum_{k}e^{i2\pi\alpha kn}\psi_{k}.$ (17)
-2. Transformation of individual terms
-Using standard trigonometric identities and Fourier algebra, one finds
+2. Momentum-space equation
+Using standard trigonometric identities, the Fourier transform yields
 $\sum_{n}e^{-i2\pi\alpha kn}\cos(2\pi\alpha n)\phi_{n\pm 1}$ $=\cos\bigl[2\pi\alpha(k\pm 1)\bigr]\psi_{k\pm 1},$ (18)
@@
 $\sum_{n}e^{-i2\pi\alpha kn}\cos(2\pi\alpha n)\phi_{n}$ $=\frac{1}{2}\bigl(\psi_{k-1}+\psi_{k+1}\bigr).$ (20)
-3. Momentum-space equation
-Substituting the above relations into the Fourier transform of Eq. (16), we arrive at
+Substituting into Eq. (16) gives
 $\cos\bigl[2\pi\alpha(k+1)\bigr]\psi_{k+1}+\cos\bigl[2\pi\alpha(k-1)\bigr]\psi_{k-1}$ (21)
 $\quad+\frac{V}{2}(\psi_{k-1}-\psi_{k+1})=\frac{E}{2}(\psi_{k-1}+\psi_{k+1}),$
-which can be rearranged into the compact form given in Eq. (15). This completes the derivation of the momentum-space representation of the QNE model.
+which reduces to Eq. (15).